OpenStudy (anonymous):

Which equation is in standard form and represents a line with slope 5 through the point (0, 2)

OpenStudy (anonymous):

i havent taken algebra 2 yet please help

OpenStudy (anonymous):

ok put (y-y1)=m(x-x1) .. putt the values of slope as m and x1 and y1 as the corrdinates of the point given (x1,y1)=(0,2) u can get ur answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(y-y1)=m(x-x1) use this equation and get ur answer mention not